# Conformal Temporal Logic Planning using Large Language Models

### v4

@@ III-E Constructing Prediction Sets using Conformal Prediction
Causal Construction of the Prediction Set: Notice that $\bar{\mathcal{C}}(\bar{\ell}_{\text{test}})$ is constructed after the entire sequence ¯$\bar{\ell}_{\text{test}}={\ell}_{\text{test}}(1),\dots,{\ell}_{\text{test}}(H_{\text{test}})$ is obtained. However, at every (test) time $t\in\{1,\dots,H_{\text{test}}\}$, the robot observes only the prompt ${\ell}_{\text{test}}(t)$ and not the whole sequence. In what follows, we construct the prediction set in a causal manner using only the current and past information. Specifically, at every time step $t$, we construct the local prediction set as [line 6-7, Alg. 2]
$\mathcal{C}({\ell}_{\text{test}}(t))=\{s~{}|~{}g(s|{\ell}_{\text{test}}(t))>1-\bar{q}\}.$ (13)
Then, the causal prediction set for ¯$\bar{\ell}_{\text{test}}$ is defined as
$\mathcal{C}(\bar{\ell}_{\text{test}})=\mathcal{C}({\ell}_{\text{test}}(1))\times\mathcal{C}({\ell}_{\text{test}}(2))\times\dots\times\mathcal{C}({\ell}_{\text{test}}(H_{\text{test}})).$ (14)
In Section IV, we show that $\bar{\mathcal{C}}(\bar{\ell}_{\text{test}})=\mathcal{C}(\bar{\ell}_{\text{test}})$.
Remark III.2 (Dataset-Conditional Guarantee)
The probabilistic guarantee in (11) is marginal in the sense that the probability is over both the sampling of the calibration set ${\mathcal{M}}$ and the validation point ¯$\bar{\ell}_{\text{test}}$. Thus, a new calibration set will be needed for every single test data point ¯$\bar{\ell}_{\text{test}}$ to ensure the desired coverage level. A dataset-conditional guarantee which holds for a fixed calibration set can also be applied [83, 41].
IV Probabilistic Task Satisfaction Guarantees
In this section, we show that given any (validation) scenario $\xi_{\text{test}}=\{{\mathcal{A}}_{\text{test}},\phi_{\text{test}},H_{\text{test}},\Omega_{\text{test}}\}$ drawn from a distribution ${\mathcal{D}}$, the probability that HERACLEs will design a plan satisfying the LTL-NL formula is at least $1-\alpha$, where $\alpha$ is the coverage level used to construct the prediction sets. To show this, we need first to show the following result. The proofs follow a similar logic to the ones in [41].
Proposition IV.1
@@ IV Probabilistic Task Satisfaction Guarantees
It suffices to show that if $\tau_{\phi}$ belongs to $\bar{{\mathcal{C}}}(\bar{\ell}_{\text{test}})$ then it also belongs to ${{\mathcal{C}}}(\bar{\ell}_{\text{test}})$ and vice-versa. First, we show that if $\tau_{\phi}\in\bar{{\mathcal{C}}}(\bar{\ell}_{\text{test}})$ then $\tau_{\phi}\in{\mathcal{C}}(\bar{\ell}_{\text{test}})$. Since $\tau_{\phi}\in\bar{\mathcal{C}}(\bar{\ell}_{\text{test}})$, then we have that $\min_{t\in\{1,\dots,H_{\text{test}}\}}g(\tau_{\phi}(t)|{\ell}_{\text{test}}(t))>1-\bar{q}$ due to (8). This means that the score $g(\tau_{\phi}(t)|{\ell}_{\text{test}}(t))>1-\bar{q}$, for all $t\in\{1,\dots,H_{\text{test}}\}$. Thus, $\tau_{\phi}(t)\in{\mathcal{C}}({\ell}_{\text{test}}(t))$, for all $t\in\{1,\dots,H_{\text{test}}\}$. By definition of ${\mathcal{C}}(\bar{\ell}_{\text{test}})$ in (14), this implies that $\tau_{\phi}\in{\mathcal{C}}(\bar{\ell}_{\text{test}})$. These steps hold in the other direction too showing that if $\tau_{\phi}\in{\mathcal{C}}(\bar{\ell}_{\text{test}})$ then $\tau_{\phi}\in\bar{{\mathcal{C}}}(\bar{\ell}_{\text{test}})$. ∎
Theorem IV.2
@@
Consider test mission scenarios $\xi_{\text{test}}$ drawn from ${\mathcal{D}}$. Assume that prediction sets are constructed causally with coverage level $1-\alpha$ and that the robot seeks help from a user whenever the local prediction set ${\mathcal{C}}({\ell}_{\text{test}}(t))$, defined in (13), is not singleton; see Section III-D. If the LTL task planning algorithm is complete and Assumption II.1 holds, then the completion rate over new test scenarios (and the randomness of the calibration sets) drawn from ${\mathcal{D}}$ is at least $1-\alpha$ [^6] Theorem IV.2 implies that HERACLEs can achieve $1-\alpha$ mission success rates even if alternative symbolic planners are used as long as they are complete.
Proof:
@@
If the symbolic planning algorithm is complete, then this means if there exists a solution it will find it. By solution, here we refer to a sequence of sub-tasks in the unknown environment that, if completed, the LTL mission will be satisfied. Since the scenarios $\xi_{\text{test}}$ are drawn from ${\mathcal{D}}$, this means that they are feasible by assumption (see Section III-E). This equivalently means that if the symbolic task planning algorithm is complete, any failures of Alg. 1 in finding a correct plan are not attributed to the symbolic planner. Under this setting, the following three cases may occur as the robot designs its plan. Case I: We have that $|\mathcal{C}({\ell}_{\text{test}}(t))|=1$, $\forall t\in\{1,\dots,H_{\text{test}}\}$ (possibly after asking for help from the symbolic planner) and $\tau_{\phi,\text{test}}\in{\mathcal{C}}(\bar{\ell}_{\text{test}})$ where $\tau_{\phi,\text{test}}$ is the ground truth plan and prediction set ${\mathcal{C}}(\bar{\ell}_{\text{test}})$ is defined as in (14). In this case, the robot will select the correct plan. Case II: We have that $|\mathcal{C}({\ell}_{\text{test}}(t))|>1$ (even after asking for help from the symbolic planner), for at least one time step $t\in\{1,\dots,H_{\text{test}}\}$ and $\tau_{\phi,\text{test}}\in{\mathcal{C}}(\bar{\ell}_{\text{test}})$. In this case, the robots will select the correct plan assuming users who faithfully provide help. Case III: We have that $\tau_{\phi,\text{test}}\notin{\mathcal{C}}(\bar{\ell}_{\text{test}})$. The latter means that there exists at least one time step $t$ such that $\tau_{\phi,\text{test}}(t)\notin{\mathcal{C}}({\ell}_{\text{test}}(t))$. In this case, the robot will compute an incorrect plan. Notice that we do not know which of these three cases will occur at each step $t$, since the ground truth plan for the test mission scenario is unknown. However, the probability that either Case I or II will occur is equivalent to the probability $P(\tau_{\phi,\text{test}}\in{\mathcal{C}}(\bar{\ell}_{\text{test}}))$. Due to Proposition IV.1 and (11), we have that $P(\tau_{\phi,\text{test}}\in{\mathcal{C}}(\bar{\ell}_{\text{test}}))\geq 1-\alpha$. Thus, either of Case I and II will occur with probability that is at least equal to $1-\alpha$. Since Cases I-III are mutually and collectively exhaustive, we conclude that the probability that Case III will occur is less than $\alpha$. This means that the probability of HERACLEs generating a correct plan is at least $1-\alpha$. Due to Assumption II.1, we also note that the mission success rate is at least $1-\alpha$ completing the proof. ∎
Remark IV.3 (Relaxing Assumption II.1)
Assumption II.1 requires perfect execution of robot skills and is required to provide the mission completion guarantees discussed in Theorem IV.2. Relaxing this assumption necessitates integrating skill imperfections in the conformal prediction component. Construction of such prediction sets requires developing score functions that (heuristically) quantify the likelihood of success of a given skill from the current state. A potential approach to address this is by leveraging the affordance functions learned in [50]. Selecting robot states that maximize these affordance functions can be accomplished by leveraging action-oriented semantic maps [84]. Subsequently, we can compute the product of this skill score function with the score function that heuristically captures LLM uncertainty (modeled by the LLM confidence value). This way, a higher product score indicates a greater likelihood that the corresponding action is correct and will be executed successfully. Using this product score function, CP can be applied as in Section III-E to construct prediction sets. In cases of non-singleton sets, among all correct actions (if any) that are included in the prediction set, the user will select the one with the highest likelihood of successful execution. Formally relaxing this assumption and designing LLM-based planners capable of reasoning about uncertainty in robot skills is part of our future work.
Remark IV.4 (Relaxing i.i.d. Assumption)
Theorem IV.2 holds under the assumption that validation and test scenarios are sampled independently from a distribution ${\mathcal{D}}$. We emphasize that this a common requirement in related works that employ CP for uncertainty quantification; see e.g., the works in [65, 60, 41, 66, 67, 68, 69, 70]. This assumption can be relaxed by employing robust CP to obtain valid prediction sets for all distributions ${\mathcal{D}}^{\prime}$ that are ‘close’ to $D$ (based on the flimit-$f-$divergence) [85]. Integrating HERACLEs with robust CP is out of the scope of this work.
V Numerical Experiments

### v5

@@ 3.5. Constructing Prediction Sets
Causal Construction of the Prediction Set: Notice that $\bar{\mathcal{C}}(\bar{\ell}_{\text{test}})$ is constructed after the entire sequence $\bar{\ell}_{\text{test}}={\ell}_{\text{test}}(1),\dots,{\ell}_{\text{test}}(H_{\text{test}})$ is obtained. However, at every (test) time $t\in\{1,\dots,H_{\text{test}}\}$, the robot observes only the prompt ${\ell}_{\text{test}}(t)$ and not the whole sequence. In what follows, we construct the prediction set in a causal manner using only the current and past information. Specifically, at every time step $t$, we construct the local prediction set as [line 6-7, Alg. 2]
(12) $\mathcal{C}({\ell}_{\text{test}}(t))=\{s~|~g(s|{\ell}_{\text{test}}(t))>1-\bar{q}\}.$
Then, the causal prediction set for $\bar{\ell}_{\text{test}}$ is defined as
(13) $\mathcal{C}(\bar{\ell}_{\text{test}})=\mathcal{C}({\ell}_{\text{test}}(1))\times\mathcal{C}({\ell}_{\text{test}}(2))\times\dots\times\mathcal{C}({\ell}_{\text{test}}(H_{\text{test}})).$
In Section 4, we show that $\bar{\mathcal{C}}(\bar{\ell}_{\text{test}})=\mathcal{C}(\bar{\ell}_{\text{test}})$.
4. Probabilistic Task Satisfaction Guarantees
In this section, we show that given any (validation) scenario $\xi_{\text{test}}=\{{\mathcal{A}}_{\text{test}},\phi_{\text{test}},H_{\text{test}},\Omega_{\text{test}}\}$ drawn from a distribution $\mathcal{D}$}, the probability that HERACLEs will design a plan satisfying the LTL-NL formula is at least $1-\alpha$, where $\alpha$ is the coverage level used to construct the prediction sets. Recall from Section 3.5 that we assume that all mission scenarios drawn from $\mathcal{D}$} admit a unique feasible plan. This means that only one sequence of sub-tasks can lead to mission completion, and there is a single robot plan that can successfully execute them. A discussion on how this assumption can be relaxed can be found in Appendix B. To show this, we need first to show the following result. The proofs follow a similar logic as in (Ren et al., 2023).
Proposition 4.1.
@@ 4. Probabilistic Task Satisfaction Guarantees
It suffices to show that if $\tau_{\phi}\in\bar{{\mathcal{C}}}(\bar{\ell}_{\text{test}})$ then $\tau_{\phi}\in{{\mathcal{C}}}(\bar{\ell}_{\text{test}})$ and vice-versa. First, we show that if $\tau_{\phi}\in\bar{{\mathcal{C}}}(\bar{\ell}_{\text{test}})$ then $\tau_{\phi}\in{\mathcal{C}}(\bar{\ell}_{\text{test}})$. Since $\tau_{\phi}\in\bar{\mathcal{C}}(\bar{\ell}_{\text{test}})$, then we have that
$\min_{t\in\{1,\dots,H_{\text{test}}\}}g(\tau_{\phi}(t)|{\ell}_{\text{test}}(t))>1-\bar{q},$
due to (7). This means that the score $g(\tau_{\phi}(t)|{\ell}_{\text{test}}(t))>1-\bar{q}$, for all $t\in\{1,\dots,H_{\text{test}}\}$. Thus, $\tau_{\phi}(t)\in{\mathcal{C}}({\ell}_{\text{test}}(t))$, for all $t\in\{1,\dots,H_{\text{test}}\}$. By definition of ${\mathcal{C}}(\bar{\ell}_{\text{test}})$ in (13) such that $\tau_{\phi}\in{\mathcal{C}}(\bar{\ell}_{\text{test}})$. These steps hold in the other direction too showing that if $\tau_{\phi}\in{\mathcal{C}}(\bar{\ell}_{\text{test}})$ then $\tau_{\phi}\in\bar{{\mathcal{C}}}(\bar{\ell}_{\text{test}})$. ∎
Theorem 4.2.
@@
Assume that all scenarios drawn from $\mathcal{D}$} have a unique feasible plan and that the prediction sets are constructed causally with coverage level $1-\alpha$. If (i) the LTL task planning algorithm is complete—that is, it can always generate an alternative sub-task, if it exists, whenever the current one is infeasible due to environmental structure, and it can generate a new sub-task as soon as the current one is completed—and (ii) Assumption 2.1 holds, then the completion rate over new test scenarios $\xi_{\text{test}}$ (and over the randomness of the calibration sets), drawn from $\mathcal{D}$}, is guaranteed to be at least $1-\alpha$.
Proof.
@@
If the symbolic planning algorithm is complete, then this means if there exists a solution it will find it. By solution, here we refer to a sequence of sub-tasks in the unknown environment that, if completed, the LTL mission will be satisfied. Since the scenarios $\xi_{\text{test}}$ are drawn from $\mathcal{D}$}, this means that they are feasible by assumption (see Section 3.5). This equivalently means that if the symbolic task planning algorithm is complete, any failures of Alg. 1 in finding a correct plan are not attributed to the symbolic planner. Under this setting, the following three cases may occur as the robot designs its plan. Case I: We have that $|\mathcal{C}({\ell}_{\text{test}}(t))|=1$, $\forall t\in\{1,\dots,H_{\text{test}}\}$ and $\tau_{\phi,\text{test}}\in{\mathcal{C}}(\bar{\ell}_{\text{test}})$ where $\tau_{\phi,\text{test}}$ is the ground truth plan and prediction set ${\mathcal{C}}(\bar{\ell}_{\text{test}})$ is defined as in (13). In this case, the robot will select the correct plan. Case II: We have that $|\mathcal{C}({\ell}_{\text{test}}(t))|>1$ for at least one time step $t\in\{1,\dots,H_{\text{test}}\}$ and $\tau_{\phi,\text{test}}\in{\mathcal{C}}(\bar{\ell}_{\text{test}})$. In this case, the robot will select the correct plan assuming users who faithfully provide help. Case III: We have that $\tau_{\phi,\text{test}}\notin{\mathcal{C}}(\bar{\ell}_{\text{test}})$. The latter means that there exists at least one time step $t$ such that $\tau_{\phi,\text{test}}(t)\notin{\mathcal{C}}({\ell}_{\text{test}}(t))$. In this case, the user is unable to provide help and instead halts the operation. Since scenarios drawn from $\mathcal{D}$} admit a unique feasible plan by assumption, the LTL planner also fails to generate alternative sub-tasks. Consequently, in Case III, the mission terminates unsuccessfully. Notice that we do not know which of these three cases will occur at each step $t$, since the ground truth plan for the test mission scenario is unknown. However, the probability that either Case I or II will occur is equivalent to the probability $P(\tau_{\phi,\text{test}}\in{\mathcal{C}}(\bar{\ell}_{\text{test}}))$. Due to Proposition 4.1 and (10), we have that $P(\tau_{\phi,\text{test}}\in{\mathcal{C}}(\bar{\ell}_{\text{test}}))\geq 1-\alpha$. Thus, either of Case I and II will occur with probability that is at least equal to $1-\alpha$. Since Cases I-III are mutually and collectively exhaustive, we conclude that the probability that Case III will occur is less than $\alpha$. This means that the probability of HERACLEs generating a correct plan is at least $1-\alpha$. Due to Assumption 2.1, we conclude that the mission success rate is at least $1-\alpha$ completing the proof.
5. Experiments